10p - 14 - p = 4p - 9 what does p equal
Answer by ikdeep(226)

10p - 14 - p = 4p - 9
on taking like items together from both sides we get...
10p - p - 4p = -9 + 14
10p - 5p = 5
5p = 5
on dividing both sides by 5 we get
p = 1
Now if you wanna verify you answer ,,,you can put the values of p in the equation and if you get LHS=RHS,,,this means that you answer is correct..